the length of a rectangle is (2x+3)cm the width is 0.5(x+3)cm the perimeter of the rectangle is 49cm find the value of x
Found 2 solutions by ewatrrr, addingup:
Answer by ewatrrr(24785)

 
Hi,
.
P = 2L + 2w
49cm = 4x + 6 + x+ 3
40cm = 5x
8cm = x
L = 19cm and w = 11/2 cm
AND CHECKING
2L + 2w = p
38cm + 11cm = 49CM CHECKS!

Answer by addingup(3677)

Perimeter = 2L+2W
2(2x+3)+2(0.5(x+3)) = 49
4x+6+x+3 = 49 add on left
5x+9 = 49
5x = 40
x = 8
-----------------
Check:
2(8)+3 = 19
0.5*(8+3) = 5.5
Perimeter:
2L+2W = 2(19)+2(5.5) = 39+10 = 49 Correct
:
J